What is the molality of 40g KBr dissolved 800 mL water

1
Expert's answer
2022-03-13T15:18:54-0400

Solution:

solute = KBr

solvent = water (H2O)


The density of water (H2O) is around approximately 1 g/mL

Density = Mass / Volume

Mass = Density × Volume

Therefore,

Mass of H2O = (1 g/mL) × (800 mL) = 800 g

Kilograms of H2O = (800 g H2O) × (1 kg / 1000 g) = 0.8 kg H2O


The molar mass of KBr is 119 g/mol

Therefore,

Moles of KBr = (40 g KBr) × (1 mol KBr / 119 g KBr) = 0.336 mol KBr


Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of KBr solution = Moles of KBr / Kilograms of H2O

Molality of KBr solution = (0.336 mol) / (0.8 kg) = 0.42 mol/kg = 0.42 m

Molality of KBr solution = 0.42 m


Answer: The molality of KBr solution is 0.42 m